Solve the system of 2 equations and 2 unknowns using the echelon method.
-4x - 5y = 15
-2x + 2y = -6
(___, ___)
Thanks.

Answer by Cromlix(4381)

Hi there,
Start with the given system of equations:
system:
-4x - 5y = 15
-2x + 2y = -6
Now you have to change the 'x' values
or the 'y' values to equal values, but
opposite signs.
Here we will multiply the lower set by -2
-4x - 5y = 15
-2x + 2y = -6 (x -2)
-4x - 5y = 15
4x - 4y = 12
Add the left parts of the equation together,
then the right parts.
(-4x - 5y) + (4x - 4y) = 15 + 12
The 'x' values vanish (-4x + 4x = 0)
-9y = 27
Divide both sides by -9
y = -3
Substitute y = -3 into
-4x - 5y = 15
-4x -5(-3) = 15
-4x = 15 - 15
-4x = 0
x = 0
{0,-3)
Hope this helps:-)